OpenStudy (anonymous):

z varies directly with x^2 and inversely with y. When x=2 and y=4, z=3. what is the value of z when x=4 and y=9?

jimthompson5910 (jim_thompson5910):

z varies directly with x^2 and inversely with y means we have this equation \[\Large z = \frac{kx^2}{y}\] where k is some fixed number

jimthompson5910 (jim_thompson5910):

Plug in x=2,y=4, z=3 and solve for k \[\Large z = \frac{kx^2}{y}\] \[\Large 3 = \frac{k(2)^2}{4}\] \[\Large 3 = \frac{4k}{4}\] \[\Large 3 = k\] \[\Large k = 3\] With me so far?